20b)Use Cauchy integral formula for derivatives evaluate ∮c Sin 2πz/(z+1)² (z²+4)dz, where C is |z + 2| <2.

Explain the problem with step by step process?


1
Expert's answer
2021-09-16T02:08:26-0400

"\u222ec z\u00b2-2z\/(z+1)^2 (z\u00b2+4)dz"

"=\u222e\ud835\udc35(\ud835\udc56\/2)1\ud835\udc672+1\ud835\udc51\ud835\udc67=\u222e"

"=∮_{B2}{i/2} (1/2i)/(z+1)^2 dz-\oint_{B2}"

"=\frac{1}{2i}-\frac{1}{2i}"

"=0"